OpenStudy (ledah):

Simplify the following: a) (2x)(3x^2)= b) (4xy^2)^3= c) (4x/3y)^2= d) 2x^4y/6xy^2=

OpenStudy (anonymous):

a= 6x^3, b= 64x^3y^6, c = 16x^2/9y^2, d= x^3/3y